OpenStudy (anonymous):

7/8u^2 divided by 56/8u

OpenStudy (lgbasallote):

hint: \[\huge \frac 7{8u^2} \div \frac{56}{8u} \implies \frac 7{8u^2} \times \frac{8u}{56}\] does that help?

OpenStudy (anonymous):

Yes thank you so much

OpenStudy (lgbasallote):

welcome